# Merge then Realign: Simple and Effective Modality-Incremental Continual Learning for Multimodal LLMs

### v1

@@ Appendix B Implementation Details
Hyperparameters Std-Stage 1 Std-Stage 2 Realigning
@@
Table 7: Hyperparameters for each training stage. Std-stage 1 and Std-stage 2 refer to the stage 1 and 2 of the standard MLLM training process.
Appendix C Evaluation Details
@@
Appendix D Implementation of Baselines
@@
  • Replay leverage a replay dataset $R_{i}\leftarrow$ sample $r\%$ data from $\{D_{1},D_{2},\dots,D_{i-1}\}$ when training on a new modality $M_{i}$. It can be seen as training on a joint dataset of $R_{i}$ and $D_{i}$.
  • EWC firstly estimates the Fisher information matrix $F_{i-1}$ of the last training stage $i-1$ as:
@@
  • PathWeave leverages Adapter-in-Adapter (AnA) modules. In our implementation, the AnA modules are injected in the LLM rather than the connector for a fair comparison. The rank of AnA is consistent with the LoRA rank of other baselines, which is 128.
Appendix E Complete Raw Data

### v2

@@ A.6 Why is the Performance Degradation More Severe in Reverse Training Order
Hyperparameters Pre-Training Fine-Tuning Realigning
@@
Table 9: Hyperparameters for each training process. Pre-Training and Fine-Tuning refer to the two stages of the standard MLLM training process.
Appendix B Dataset Details
Appendix C Experimental Details
C.1 Implementation Details
  • Audio: We use $\mathrm{BEATs_{iter3+}(AS2M)}$ (Chen et al., 2023b) as the pre-trained audio encoder, a Q-Former (Li et al., 2023) initialized from the pre-trained bert-base-uncased (Kenton and Toutanova, 2019) as the connector. The number of query tokens is set to 32.
  • Point Cloud: We use Point-BERT-v1.2 (Xu et al., 2024) as the pre-trained point cloud encoder, a randomly initialized MLP as the connector.
We set the hyperparameters mainly following previous works (Liu et al., 2024b, a), as listed in Table 9. For training that involves updating the LLM backbone, we utilize parameter-efficient fine-tuning with LoRA (Hu et al., 2022) applied across all linear modules within the LLM, setting the LoRA rank to 128 and the alpha parameter to 128. All the experiments are conducted on a single NVIDIA RTX A6000 48G with FP16.
@@ C.1 Implementation Details
C.2 Evaluation Details
C.3 Implementation of Baselines
@@
  • Replay leverage a small replay dataset $R_{i}\leftarrow$ randomly sample rr% data from $\{D_{1},D_{2},\dots,D_{i-1}\}$. When training on a new modality $M_{i}$, it trains on the joint dataset of $R_{i}$ and $D_{i}$.
  • EWC firstly estimates the Fisher information matrix $F_{i-1}$ of the last training stage i1i-1 as:
@@
  • PathWeave leverages Adapter-in-Adapter (AnA) modules. In our implementation, the AnA modules are injected into the LLM rather than the connector for a fair comparison. The rank of AnA is consistent with the LoRA rank of other baselines, which is 128. In their original settings, PathWeave removes the newly added modules when testing the former modalities. However, this results in the inability to perform cross-modality tasks, which are common in real applications. Therefore, we do not remove them for a fair comparison.
For each baseline, all the common parameters about training MLLM itself are the same and set as default in Table 9.
Appendix D Complete Raw Data